OpenStudy (anonymous):

simplify: sqrt60 x sqrt105 (x=multiply)

OpenStudy (anonymous):

\[60=2^2\times 3\times 5\] \[105=3\times 5\times 7\] \[60\times 105=2^2\times 3^2\times 5^2\times 7\] \[\sqrt{60\times 105}=\sqrt{2^2\times 3^2\times 5^2\times 7}=2\times 3\times 5\times \sqrt{7}=30\times \sqrt{7}\]

OpenStudy (anonymous):

\[\sqrt{15*4}*\sqrt{105}\] \[2\sqrt{15}*\sqrt{15*7}\] \[2*15\sqrt{7}\] \[30\sqrt{7}\]

OpenStudy (anonymous):

ok i get it now thx